Balanced equation 2NaOH (aq) + CO2 --> Na2CO3 (aq) + H2O Calculate the number of grams of carbon dioxide that can be absorbed by complete reaction with a solution that contains 5.00 g of sodium hydroxide.
I need to know how to do the equation as well. I was never taught how to do this.
1
Expert's answer
2013-01-23T11:16:40-0500
2NaOH + CO2 = Na2CO3 + H2O
First calculate the amount of substance of sodiumhydroxide:
n(NaOH) = m/M
n(NaOH) = 5/M
M(NaOH) = 23 + 16 + 1 = 40 g/mol
n(NaOH) = 5 g / 40 g/mol = 0.125 g/mol
n(CO2) = 1/2 n (NaOH)
n(CO2) = 0.125/2 = 0.0625 mol
m(CO2) = M*n
M(CO2) = 12 + 16*2 = 44 g/mol
m(CO2) = 44 g/mol * 0.0625 mol = 2.75 g
Answer: 2.75 g